# The simplest model of a scalarized black hole in the Einstein-Klein-Gordon theory

### v1

@@ Abstract
We investigate scalarized black holes in the Einstein-minimally coupled scalar theory with a negative potential $V(\phi)=-\alpha^{2}\phi^{6}$. The tachyonic instability is absent from analyzing the linearized scalar equation, which could not allow for spontaneous scalarization. However, we obtain the black hole solutions with scalar hair by solving three full equations because this scalar potential violates the weak energy condition. This shows clearly that scalarized black holes can be obtained without introducing a non-minimal scalar coupling term. We perform the stability analysis for scalarized black holes by adopting radial perturbations, implying that all scalarized black holes belonging to a single branch are unstable.
@@ 1 Introduction
Recently, it is well known that infinite branches of scalarized (charged) black holes were obtained numerically from the Einstein-Gauss-Bonnet-scalar theory [7, 8, 9, 10] (Einstein-Maxwell-scalar theory [11]) through the non-minimal coupling of a scalar function to the Gauss-Bonnet term (Maxwell term). In these linearized theories, the tachyonic instability of $s(l=0)$-scalar mode propagating around the GR black holes indicates the onset of spontaneous scalarization. This arose from either the negative Gauss-Bonnet coupling term ($-48\lambda^{2}M^{2}/r^{6}$) or the negative Maxwell coupling term($-2\alpha Q^{2}/r^{4}$) where $\lambda^{2}$ and $\alpha$ are positive scalar coupling parameters. These negative coupling terms induce actually potential wells near the horizon and as coupling parameters increase, leading to tachyonic instabilities. Furthermore, the linearized static scalar equation has played an important role in obtaining scalar clouds (bound states). Requiring an asymptotically vanishing scalar, the condition for obtaining a smooth scalar selects a discrete set of the bifurcation points for scalarized solutions: $M/\lambda=\{0.587,0.226,0.140,\cdots\}~{}(\alpha(q=0.7)=\{8.019,40.84,99.89,\cdots\})$ for the Gauss-Bonnet term [12] (as well as the Maxwell term [13]). They have admitted scalar hairy black holes in the full theory. Actually, it has induced infinite branches of scalarized [charged] black holes: $n=0(0<M/\lambda<0.587),1(0<M/\lambda<0.226),2(0<M/\lambda<0.140),\cdots[n=0(\alpha>8.019),n=1(\alpha>40.84),n=2(\alpha>99.89),\cdots]$.
@@
In this work, we focus on introducing a newly simple mechanism to obtain scalarized black holes from a negative potential term $V(\phi)=-\alpha^{2}\phi^{6}<0$ in the Einstein-minimally coupled scalar theory without depending on any scalar coupling terms. Hence, the tachyonic instability is absent. In this case, we obtain a single branch of scalarized black holes in Einstein-minimally coupled scalar theory. Carrying out stability analysis for scalarized black holes by adopting radial perturbations, we find that all scalarized black holes belonging to a single branch are unstable. This implies that even though one obtains easily scalarized black hole from the simplest scalar-tensor action, they are hard to survive for further implications.
@@
The present work is inspired from the hairy black hole solutions in Einstein-Weyl-massive conformally coupled scalar theory where the tachyonic scalar mass has played an important role in obtaining scalar hairy black holes [30]. In the absence of tachyonic mass term ($-m^{2}\phi^{2}$), we have obtained the non-BBMB black hole solution in the new massive conformal gravity [31]. Also, this work is motivated strongly from finding scalarized black hole in the Einstein-minimally coupled scalar theory with an asymmetric scalar potential which contains a negative region, indicating a violation of the weak energy condition (WEC) [32, 33, 34, 35]. Besides, the asymmetric potential has also been employed to construct the fermionic stars [36, 37]. A symmetric potential of $V(\phi)=\mu\phi^{2}-\Lambda\phi^{4}$ violating the WEC was also introduced to obtain the scalarized black holes [38]. The no-hair theorem suggests that asymptotically flat black holes with scalar hair do not exist if a scalar matter satisfies the WEC [39]. Hence, if the WEC for a scalar matter is violated, scalarized black holes could be found from the Einstein-minimally coupled scalar theory. As far as we know, our model corresponds to a simplest model which violates the WEC because of a negative potential and thus, it could induce scalarized black holes without introducing a non-minimal scalar coupling term. Other types of scalarized black holes in the Einstein-minimally coupled scalar theory can be found in Ref. [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. Besides, recently this similar concept has been adopted to construct a traversable wormhole in the Einstein-3-form theory with the Higgs-like potential, which is sufficient to violate the null energy condition where the phantom field is no longer needed and the kinetic term still can remain the correct sign [51].
@@ 2 No tachyonic instability of Schwarzschild black hole
$S_{\rm Es}=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\Big{[}\frac{R}{16\pi G}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-V(\phi)\Big{]},$ (1)
@@ 3 Scalarized black holes
with $N(r)=1-2m(r)/r$ where $m(r)$ denotes the Misner-Sharp mass function. We note that $m(\infty)=M$, the total mass of the configuration. Hereafter, we choose $\alpha=1$ for simplicity.
@@
Finally, introducing three dimensionless parameters
@@ 5 Discussions
We have explored scalarized black holes in the Einstein-minimally coupled scalar theory with a negative potential $V(\phi)=-\alpha^{2}\phi^{6}$. The tachyonic instability is absent from analyzing the linearized scalar equation. In this case, one could not meet a condition for spontaneous scalarization and thus, the linearized stability provided no further information on scalarization.
However, this negative scalar potential with $\alpha=1$ violates the weak energy condition (WEC). It is well-known that if the WEC for a scalar matter is violated, scalarized black holes could be found from the Einstein-minimally coupled scalar theory without introducing a non-minimal scalar coupling to matter. Thus, we have obtained the black hole solutions with scalar hair by solving three nonlinear equations. It includes a single branch of scalarized black holes only because tachyonic intability is absent. Different scalarized black holes are encoded in different $\phi_{H}$ because we have chosen $\alpha=1$. Furthermore, we have studied their thermodynamic properties by introducing reduced horizon area and Hawking temperature.
@@
Finally, it would be interesting to obtain the other solutions of scalarized black holes by choosing a more simpler form of the potential, for instance $V(\phi)=-m^{2}_{\phi}\phi^{2},-\Lambda\phi^{4}$ with $m^{2}_{\phi}$ and $\Lambda$ are positive constants, since the previous analysis of Eqs.(8) and (9) in [38] doesn’t rule out the possible existence of scalarized black hole solutions. Therefore, it is worth to investigate such possibilities and report them in the future.

### v3

# Simplest model of a scalarized black hole in the Einstein-Klein-Gordon theory

@@ Abstract
We investigate scalarized black holes in the Einstein-minimally coupled scalar theory with a negative potential $V(\phi)=-\alpha^{2}\phi^{6}$. The tachyonic instability is absent from analyzing the linearized scalar equation, which could not allow for spontaneous scalarization. However, we obtain the black hole solutions with scalar hair by solving three full equations because this scalar potential violates the weak energy condition. This shows clearly that a single branch of scalarized black holes can be obtained without introducing a non-minimal scalar coupling term. We perform the stability analysis for scalarized black holes by adopting radial perturbations, implying that all scalarized black holes belonging to a single branch are unstable.
@@ 1 Introduction
Recently, the spontaneous scalarization implies that infinite branches of scalarized (charged) black holes were obtained numerically from the Einstein-Gauss-Bonnet-scalar theory [7, 8, 9, 10] (Einstein-Maxwell-scalar theory [11]) through the non-minimal coupling function $f(\phi)$ with $f(0)=0,f^{\prime}(0)=0,f^{\prime\prime}(0)\not=0$ to the Gauss-Bonnet term (Maxwell term). In these linearized theories (¯$\bar{\nabla}^{2}\delta\phi-m^{2}_{\rm eff}\delta\phi=0$), the tachyonic instability of $s(l=0)$-scalar mode propagating around the GR black holes indicates the onset of spontaneous scalarization. This arose from either the negative Gauss-Bonnet coupling term ($m^{2}_{\rm eff}=-48\lambda^{2}M^{2}/r^{6}$) or the negative Maxwell coupling term ($m^{2}_{\rm eff}=-2\alpha Q^{2}/r^{4}$) where $\lambda^{2}$ and $\alpha$ are positive scalar coupling parameters. These negative coupling terms induce actually potential wells near the horizon and as coupling parameters increase, leading to tachyonic instabilities. Furthermore, the linearized static scalar equation has played an important role in obtaining scalar clouds (bound states). Requiring an asymptotically vanishing scalar, the condition for obtaining a smooth scalar selects a discrete set of the bifurcation points for scalarized solutions: $M/\lambda=\{0.587,0.226,0.140,\cdots\}~{}(\alpha(q=0.7)=\{8.019,40.84,99.89,\cdots\})$ for the Gauss-Bonnet term [12] (as well as the Maxwell term [13]). They have admitted scalar hairy black holes in the full theory. Actually, it has induced infinite branches of scalarized [charged] black holes: $n=0(0<M/\lambda<0.587),1(0<M/\lambda<0.226),2(0<M/\lambda<0.140),\cdots[n=0(\alpha>8.019),n=1(\alpha>40.84),n=2(\alpha>99.89),\cdots]$.
@@
On the other hand, the nonlinear mechanism was introduced to obtain a single branch of scalarized black holes in Einstein-scalar-Gauss-Bonnet gravity, which is surely beyond the previous spontaneous scalarization. Introducing a different coupling function $f(\phi)$ satisfying $f(0)=0,f^{\prime}(0)=0,f^{\prime\prime}(0)=0$, one found from its linearized equation (¯$\bar{\nabla}^{2}\delta\phi=0$) that Schwarzschild black hole is linearly stable against scalar perturbation, whereas it is unstable against nonlinear scalar perturbation when the amplitude of a perturbed scalar is large enough [21, 30, 31, 32]. This provides another mechanism to obtain a single branch of scalarized black holes via nonlinear scalarization [33, 34].
@@
In this work, we wish to introduce another model to find a single branch of scalarized black holes. Its linearized theory is stable against scalar perturbation, whereas its full theory indicates a violation of the weak energy condition (WEC) because the scalar potential includes negative region. We are interested in a negative potential term $V(\phi)=-\alpha^{2}\phi^{6}<0$ violating the WEC in the Einstein-minimally coupled scalar theory without introducing any coupling function $f(\phi)$. Hence, the tachyonic instability is absent. In this case, we may obtain a single branch of scalarized black holes in Einstein-minimally coupled scalar theory. Carrying out stability analysis for scalarized black holes by adopting radial perturbations, we find that all scalarized black holes belonging to a single branch are unstable. This implies that even though one obtains easily scalarized black hole from the simplest scalar-tensor action, they are hard to survive for further implications.
@@
The present work is motivated partly from the hairy black hole solutions in Einstein-Weyl-massive conformally coupled scalar theory where the tachyonic scalar mass has played an important role in obtaining scalar hairy black holes [35]. In this case, in the absence of tachyonic mass term, we have obtained the non-BBMB black hole solution in the new massive conformal gravity [36]. Also, this work is motivated partly from finding scalarized black hole in the Einstein-minimally coupled scalar theory with an asymmetric scalar potential which contains a negative region, indicating a violation of the WEC [37, 38, 39, 40]. Besides, the asymmetric potential has also been employed to construct the fermionic stars [41, 42]. A symmetric potential violating the WEC was also introduced to obtain the scalarized black holes [43], which can be smoothly connected with the counterpart gravitating scalaron in the small horizon limit [44]. Note that this gravitating scalaron can possess the positive Arnowitt-Deser-Misner (ADM) mass but previously another type of gravitating scalaron with negative ADM mass was constructed by employing the Higgs-like potential with a phantom field [45], hence this demonstrates that the use of phantom field can be avoided for the construction of gravitating scalaron. The no-hair theorem suggests that asymptotically flat black holes with scalar hair do not exist if a scalar matter satisfies the WEC [46]. Hence, if the WEC for a scalar matter is violated, scalarized black holes could be found from the Einstein-minimally coupled scalar theory. As far as we know, our model corresponds to a simplest model which violates the WEC because of a negative potential and thus, it could induce scalarized black holes without introducing a non-minimal scalar coupling term. Other types of scalarized black holes in the Einstein-minimally coupled scalar theory can be found in Ref. [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Besides, recently this similar concept has been adopted to construct a traversable wormhole in the Einstein-3-form theory with the Higgs-like potential, which is sufficient to violate the null energy condition where the phantom field is no longer needed and the kinetic term still can remain the correct sign [59].
@@ 2 No tachyonic instability of Schwarzschild black hole
$S_{\rm Es}=\int d^{4}x\sqrt{-g}\Big{[}\frac{R}{16\pi G}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-V(\phi)\Big{]},$ (1)
@@ 3 Scalarized black holes
with $N(r)=1-2m(r)/r$ where $m(r)$ denotes the Misner-Sharp mass function. We note that $m(\infty)=M$, the total mass of the configuration.
@@
Finally, introducing two dimensionless parameters
@@ 5 Discussions
It is clear that the tachyonic instability as the onset of spontaneous scalarization indicates infinite branches of scalarized black holes. We have explored a single branch of scalarized black holes in the Einstein-minimally coupled scalar theory with a negative potential $V(\phi)=-\alpha^{2}\phi^{6}$. Here, the tachyonic instability is absent and thus, it plays no role in predicting infinite branches of scalarized black holes. In this case, one could not meet a condition for spontaneous scalarization, but one meets a condition for nonlinear scalarization to obtain a single branch of scalarized black holes in the Einstein-Gauss-Bonnet-scalar theory with a coupling function $f(\phi)$. This suggests that tachyonic instability is a necessary condition to obtain infinite branches of scalarized black holes.
To generate a single branch of scalarized black hole, there are some sources of nonlinear instability [21, 30, 31, 32], conformal scalar coupling [5, 6, 14], superradiant instability [65], and violation of the WEC [37, 38, 39, 40]. In our work, it is important to note that the negative scalar potential with $\alpha=1$ violates the weak energy condition (WEC). It is well-known that if the WEC for a scalar matter is violated, scalarized black holes could be found from the Einstein-minimally coupled scalar theory without introducing any scalar coupling function $f(\phi)$ to matter. Thus, we have obtained the black hole solutions with scalar hair by solving three nonlinear equations. It includes a single branch of scalarized black holes only because tachyonic intability is absent. Different scalarized black holes are encoded in different $\phi_{H}$ because we have chosen $\alpha=1$. Furthermore, we have studied their thermodynamic properties by introducing reduced horizon area and Hawking temperature.
@@
Finally, it would be interesting to obtain the other solutions of scalarized black holes by choosing a more simpler form of the potential, for instance $V(\phi)=-m^{2}_{\phi}\phi^{2},-\Lambda\phi^{4}$ with $m^{2}_{\phi}$ and $\Lambda$ are positive constants, since the previous analysis of Eqs.(8) and (9) in [43] doesn’t rule out the possible existence of scalarized black hole solutions. Therefore, it is meaningful to investigate such possibilities and report them in the future.